# HLStrans: Dataset for LLM-Driven C-to-HLS Hardware Code Synthesis

### v1

@@ Abstract
High-level synthesis (HLS) enables software developers to describe and implement hardware at a higher level of abstraction by using C/C++ instead of traditional hardware description languages to automatically generate FPGA-ready designs. However, generating HLS code significantly differs from standard C/C++: it disallows certain coding idioms, relies on specialized libraries, and critically requires fine-grained transformations and the insertion of optimization directives (pragmas) to achieve high performance. Large language models (LLMs) have shown promise in automating such transformations, yet existing open-source datasets lack sufficient complexity and optimization diversity. To address this gap, we introduce the HLStrans dataset, a comprehensive collection of 137 distinct real word programs, each annotated with a variety of C-to-HLS transformations that yield over 23K labeled design variants. These include a broad spectrum of pragmas and code-level optimizations. We benchmark state-of-the-art LLMs on this dataset to evaluate their ability to generate synthesizable, high-performance HLS code. As part of an ongoing effort, we plan to expand the HLStrans dataset in both scale and program variety, further empowering research at the intersection of AI and hardware synthesis.
@@ 1 Introduction
Specialized computing systems, also known as domain-specific architectures tailored to particular application domains, have seen dramatic growth as workloads like machine learning, signal processing, and data analytics outstrip the capabilities of general-purpose processors. By offloading compute-intensive kernels to accelerators such as Graphics Processing Units (GPUs), Field-Programmable Gate Arrays (FPGAs), or Application-Specific Integrated Circuits (ASICs), designers achieve orders-of-magnitude improvements in throughput, latency, and energy efficiency. To capitalize on this trend, begin by profiling your target workload to identify its most critical functions, then select the most appropriate hardware substrate and toolchain to implement those kernels with maximum parallelism and minimal overhead.
High-Level Synthesis (HLS) is a methodology that elevates hardware design to higher levels of abstraction by allowing engineers to describe functionality in software languages like C or C++, from which synthesizable hardware descriptions are automatically generated. The HLS tool analyzes the code, infers parallel execution paths, inserts pipeline registers, and maps data accesses to on-chip memories, producing Register-Transfer Level (RTL) descriptions in Hardware Description Languages (HDLs) such as Verilog or VHDL. Using HLS can accelerate the hardware design process and enable software engineers to engage in hardware development.
However, HLS departs radically from conventional software development by requiring developers to adopt a narrow, hardware-centric subset of C++ and acquire domain-specific expertise. Typical HLS workflows forbid features like dynamic memory allocation, unrestricted pointer arithmetic, recursion, and most of the standard library, instead demanding code that maps directly to efficient RTL ye2024hida sohrabizadeh2022autodse . To achieve high-quality output, engineers must refactor their designs so that loops have compile-time bounds, arrays possess fixed sizes, and control logic forms strictly nested or single-producer/single-consumer structures sudvarg2024hls martin2009high ye2022scalehls . On top of this, HLS tools expose a suite of vendor-specific pragma directives for loop pipelining, unrolling, array partitioning, and dataflow streaming that developers weave into their code. Many also integrate specialized C++ libraries offering arbitrary-precision fixed-point types, hardware-friendly math routines, FIFO streams, and task-oriented dataflow constructs to bridge the gap between software abstractions and synthesizable hardware modules.
LLMs offer hope to aid designers in writing HLS code. Rather than just producing HLS code from natural language with LLMs, more recent works focus on converting existing C kernels into optimized HLS implementations collini2024c2hlsc xiong2024hlspilot bhattacharyya2024llm xu2024automated prakriya2025lift , leveraging structured inputs to yield high-performance designs rather than naive, resource-heavy translations. Despite these advances, integrating LLMs into HLS toolchains remains challenging and relatively unexplored, in large part because there is no comprehensive, open-source corpus of diverse, high-quality HLS kernels. This gap prevents practitioners and researchers from fully exploiting the code-generation and optimization capabilities of modern LLMs within their design workflows.
To address the scarcity of high-quality open-source HLS benchmarks, we introduce the HLStrans dataset, a comprehensive suite of 23K HLS kernels harvested from academic repositories. Each entry pairs a naive, functionally equivalent C implementation with multiple performance-tuned HLS variants. To assemble this corpus, we developed an automated augmentation pipeline that leverages Monte Carlo tree search (MCTS) guided by electronic design automation (EDA) tool feedback and design space exploration (DSE) tools, ensuring diverse yet compiler-friendly code transformations. Every kernel is annotated with cycle-accurate latency and detailed resource-utilization metrics obtained via synthesis, creating a rich training signal for both functional correctness and performance optimization. Building on this resource, we further fine-tune large language models using a novel reinforcement-style scheme of "performance tags" and "resource tags" (rated 1–10) that explicitly reward efficient hardware design choices.
In summary, we have the following contributions:
  • HLStrans Dataset: We introduce HLStrans, the first large-scale dataset of over 23K paired C and HLS programs for LLM-powered HLS code synthesis from software.
  • LLM-Driven HLS Data Augmentation: We propose the first effective augmentation framework that leverages LLMs, MCTS, EDA feedback, and DSE tools to generate high-quality synthetic HLS code.
  • Benchmark-Driven LLM Evaluation: Enabled by our benchmark suite, we evaluate various prompts and fine-tuning methods and demonstrate that while off-the-shelf code LLMs underperform on HLS tasks, retrieval augmentation and task-specific fine-tuning on HLStrans yield substantial improvements.
2 Background
Generating high-performance HLS code from standard C for FPGAs requires a series of systematic transformations to expose parallelism, optimize data movement, and conform to HLS-friendly coding styles. While the detailed transformations are in Appendix A.1, these transformations fall into five broad categories overall:
(a) Code restructuring: loop tiling examples
(b) Directive insertion: unroll and array partition
(c) Data-type adaptation: float to fixed_type
(d) Function transformation: sqrt example
Figure 1: HLS code optimization examples
(1) HLS Code Optimization: Code Restructuring. Refactor algorithms to expose pipelining and dataflow, apply loop tiling, memory coalescing, ping-pong buffering, and reorganize control logic for parallel or streaming execution. Figure 1(a) shows an example of loop tiling.
(2) HLS Code Optimization: Directive (Pragma) Insertion. Place HLS pragmas to guide the tool scheduler, such as data flow, pipeline, loop partition, and interface specifications, to fine-tune performance and resource usage. Figure 1(b) shows an example of pragma use for unrolling and array partitioning.
(3) HLS Code Optimization: Data-Type Adaptation. Replace generic C types with HLS-specific arbitrary-precision types: convert floating point to fixed point (ap_fixed) for resource optimization, standard integers to bit-accurate (ap_int/ap_uint), and customize bit widths to match application precision requirements. Figure 1(c) shows an example of transforming float to fixed types.
(4) HLS Code Optimization: Transformation of Functions. Transform standard C functions into HLS-optimized kernels or intrinsics (such as converting the std::sqrt function to the hls::sqrt function) to better leverage FPGA fabric and specialized accelerators. Figure 1(d) shows an example of transforming the standard sqrt function to the sqrt function of the HLS library.
(5) HLS Code Repair: HLS-Compliant Coding Style. Eliminate unsupported C constructs such as dynamic memory allocation (malloc/free), recursion, and certain pointer arithmetic patterns; restructure code to use static arrays, simple loops, and explicit handshaking for communication. Figure 2 shows an example of removing recursion, which cannot be synthesized successfully.
Figure 2: HLS Code repair example: recursion elimination
3 Related works
@@
There is an increasing body of literature on applying LLMs to generate HLS design from original C code. C2HLS collini2024c2hlsc evaluates the basic task of translating naive C++ into synthesizable HLS C++. bhattacharyya2024llm demonstrates that LLMs can automate HLS pragmas and optimizations to produce synthesizable, high-performance RTL from C on image-processing benchmarks. LIFT prakriya2025lift introduces LLM–GNN hybrid methods that automatically optimize C/C++ designs to boost FPGA performance. xu2024automated presents an LLM-driven HLS program-repair framework that transforms C/C++ into synthesizable HLS-C. HLSPilot xiong2024hlspilot extends this approach with retrieval-augmented generation and chain-of-thought prompting to deliver optimized HLS implementations across nine applications. However, to date, no work has established a large-scale dataset for thoroughly evaluating an LLM’s HLS capabilities.
First, prior HLS datasets have primarily focused on resource estimation rather than on generating HLS code from C, and available program sources remain scarce. HLSsyn bai2023towards concentrates on encompassing a diverse range of optimization pragmas into only 42 kernels for training and evaluation of design quality prediction models. HLSDataset wei2023hlsdataset , including 34 data sources and 18K samples, targets power estimates, resource, and timing estimation. MLSBench goswami2022mlsbench is an open-source dataset created with the Xilinx Vivado HLS flow, covering 17 C/C++ and 13 SystemC benchmarks drawn from Polybench, CHStone, and MachSuite. It includes only HLS log files and reports. Spector gautier2016spector comprises a collection of OpenCL applications that the authors process with the Intel OpenCL SDK to produce 8,300 distinct hardware implementations for Intel FPGAs. Dai et al. dai2018fast have open-sourced a dataset of about 1,300 designs generated by using applications from suites such as CHStone, Machsuite, and Rosetta. Second, existing datasets inadequately capture comprehensive C-to-HLS transformations, focusing largely on pragma insertion. Third, they are not organized as paired C-and-HLS examples and omit the corresponding testbenches needed for LLM-based HLS code optimization, which are not ready for LLM.
@@
Table 1: Comparison of different HLS dataset
Work Samples Sources Platform & Abstraction level Supported transformation Use Case in ML Ready for LLM
Daidai2018fast 1,300 65 FPGA HLS (2) Quality of Results Estimation No
MLSBenchgoswami2022mlsbench 6,000 30 FPGA HLS (2) No
Spectorgautier2016spector 8,300 9 FPGA HLS (2) No
HLSdatasetwei2023hlsdataset 18,876 34 FPGA HLS (1) (2) (3) (4) Quality of Results Estimation. No
HLSsynbai2023towards 42,000 42 FPGA HLS (2) Quality of Results Estimation. No
Ours 23,000 137 FPGA HLS (1)(2)(3)(4)(5) Code generation. Yes
4 High-Level Synthesis (HLS) Datasets Construction
Open-source HLS datasets are scarce and unstructured, making them insufficient for training LLMs. To address this, we will develop an automated pipeline that generates high-quality HLS datasets from the limited available resources. The datasets are released under the MG0-2.0 license (duan2024modelgo, ). [^1] https://www.modelgo.li/. To construct a high-quality dataset, we need two steps. The first step is to collect high-quality open-source HLS examples. However, human-optimized code examples are not enough. The second step is the effective data augmentation for human-optimized codes to produce effective codes. Figure 3 shows our dataset construction process.
@@ 4.1 Dataset Collection
Manually optimized codebases often suffer from inconsistent formatting and minimal documentation, making them ill-suited for LLM-driven code generation. Public kernels frequently rely on unexpanded macros and include extraneous utility functions that obscure core algorithms. Our goal is to prepare C-to-HLS designs for seamless LLM consumption by packaging each design with the following files:
@@
  • Self-contained C++ testbench $tb$ includes all test cases and validation logic necessary to verify the kernel’s outputs against expected results. We manually write the testbenches and adjust the optimized HLS code to ensure it passes all tests.
Therefore, we construct pairs $(x,y,tb)$. If the original file $x$ is synthesisable, the execution cycles from synthesis reports of $y$ must be less than $y$. If the original file $x$ is not synthesisable, $y$ should be synthesisable. Table 2 shows our HLS dataset collection including common heterogeneous benchmarks such as polybench, and other FPGA benchmarks.
Table 2: Dataset collection summary
Data Source Applications Average Line of codes Supported transformations
Polybenchpouchet2012polyhedral 28 23 (1)(2)(3)
Rodiniacong2018understanding 15 61 (1)(2)(3)
Machsuitereagen2014machsuite 16 59 (2)(3)
Rosettazhou2018rosetta 3 127 (1)(2)(3)(4)
C2HLScollini2024c2hlsc 14 47 (2)(4)(5)
PP4FPGAkastner2018parallel 4 36 (1)(2)(3)(4)
Forgebenchwanna2025forgebench 6 112 (4)
HLSfactoryabi2024hlsfactory 11 107 (1)(2)(3)(4)
Others (GitHub) 41 101 (1)(2)(3)(4)(5)
Totals 137 93 (1)(2)(3)(4)(5)
@@ 4.2 Dataset Augmentation
Given the performance, resource utilization, and compilation feedback provided by EDA toolchains, we employ Monte Carlo tree search (MCTS) and design space exploration (DSE) tools to efficiently explore the large and combinatorial action space as shown in Figure 3.
First, MCTS performs structured exploration by balancing the exploitation of high-reward actions with the exploration of uncertain or less-visited regions of the search space. The optimization policy is generated by the retrieval augmentation generation (RAG) module. The search is guided to choose the suitable policy by both the verification pipeline and a reward model. The reward model incorporates detailed feedback from the HLS toolchain, including synthesis success or failure, compile warnings, and performance metrics such as resource usage, latency, and throughput. This heuristic-driven strategy enables the agent to iteratively refine transformation sequences and produce more high-quality, synthesizable HLS designs.
Second, HLS directive design space exploration using genetic algorithms ferikoglou2023collectivehls is adopted. It inserts pipeline, unroll, and partition pragmas to produce more effective data samples. Through iterative refinement, the framework converges toward optimized and synthesizable HLS code. After collecting these C-to-HLS code pairs, we use an LLM to produce a chain-of-thought narrative that analyzes the performance and resource bottlenecks, explains why the original code cannot be synthesized, and highlights the key optimizations applied in the HLS version. This explanatory reasoning is then integrated alongside the optimized HLS output.
Figure 3: HLStrans Dataset Augmentation Framework

### v3

# HLStrans: Dataset for C-to-HLS Hardware Code Synthesis

@@ Abstract
High-Level Synthesis (HLS) enables hardware design from C/C++ kernels but requires extensive transformations, such as restructuring code, inserting pragmas, adapting data types, and repairing non-synthesizable constructs, to achieve efficient FPGA implementations. While large language models (LLMs) show promise in automating these transformations, progress has been limited by the absence of large-scale, well-structured datasets. Existing HLS datasets focus primarily on resource estimation, lack paired C and HLS examples with testbenches, and cover only a narrow set of optimizations. We introduce HLStrans, the first benchmark-scale dataset for LLM-driven C-to-HLS synthesis. HLStrans contains over 124K paired C and HLS programs for real-world applications, with full testbenches and synthesis-based annotations of latency and resource usage. The dataset systematically captures five categories of transformations and is enriched by an automated augmentation pipeline combining LLMs, Monte Carlo Tree Search (MCTS), and Design Space Exploration (DSE). We benchmark state-of-the-art LLMs on HLStrans, demonstrating that retrieval and fine-tuning significantly improve success rates and performance.
@@ 1 Introduction
Specialized computing systems, particularly FPGAs, are increasingly deployed to accelerate compute-intensive workloads in domains such as machine learning, signal processing, and data analytics. High-Level Synthesis (HLS) has emerged as a key methodology for bridging software and hardware, allowing engineers to describe functionality in C/C++ and automatically generate hardware-ready RTL. However, generating high-performance HLS code is far from a direct translation: it requires structural code refactoring, insertion of optimization pragmas, adaptation of data types, replacement of functions with hardware-friendly intrinsics, and strict compliance with HLS coding styles. Therefore, We define the C-to-HLS transformation task as follows: given a sequential C/C++ kernel, generate a synthesizable HLS implementation that achieves efficient hardware acceleration on an FPGA platform. This task exemplifies the challenges at the intersection of AI and EDA, demanding not only correctness but also hardware-aware optimization.
Recent work has demonstrated the potential of large language models (LLMs) for HLS code generation. Early studies explored direct translation from C++ to synthesizable HLS code, while others focused on automating pragma insertion, repairing unsynthesizable constructs, or leveraging retrieval-augmented and chain-of-thought prompting to improve optimization quality (Collini et al., 2024; Xiong et al., 2024; Bhattacharyya et al., 2024; Xu et al., 2024; Prakriya et al., 2025). While promising, these approaches are constrained by the lack of comprehensive benchmarks: existing evaluations are conducted on small, fragmented collections of kernels, making it difficult to reproduce results or compare methods fairly. Without a unified, large-scale dataset, it remains challenging to systematically assess or advance LLMs on the C-to-HLS task.
Although several datasets for HLS exist, such as HLSsyn (Bai et al., 2023), HLSDataset (Wei et al., 2023), MLSBench (Goswami et al., 2022), and HLSfactory (Abi-Karam et al., 2024), they fall short for this purpose. Most are designed for resource estimation rather than code transformation, are limited in scale (typically a few hundred to a few thousand kernels), and rarely include paired examples of original C code, optimized HLS code, and testbenches. Moreover, they capture only a narrow slice of transformation diversity, focusing mainly on pragma insertion and overlooking critical steps such as code restructuring, data type adaptation, and repair of unsupported C constructs. As a result, current datasets cannot serve as a benchmark foundation for training or evaluating LLMs on realistic C-to-HLS synthesis.
To address these gaps, we present HLStrans, the first benchmark-scale dataset explicitly designed for LLM-driven C-to-HLS transformation. HLStrans contains over 124,200 C and HLS pairs drawn from diverse real-world applications, covering domains such as linear algebra, machine learning, DSP, image processing, and cryptography. Each entry includes a triple: the original C kernel, an optimized HLS implementation, and a validation testbench, with annotations of latency and resource metrics obtained via synthesis. The dataset systematically captures five categories of transformations: code restructuring, pragma insertion, data type adaptation, function replacement, and HLS-compliant repair, ensuring broad coverage of hardware-oriented optimizations. To further enrich this corpus, we introduce an automated augmentation framework that combines LLMs, Monte Carlo Tree Search (MCTS), and Design Space Exploration (DSE) to generate diverse, synthesizable variants guided by synthesis feedback.
In summary, our contributions are threefold:
  • We release HLStrans, the first large-scale dataset for C-to-HLS transformation, enabling LLM training and fair benchmarking;
  • We propose a novel augmentation pipeline that produces diverse, high-quality HLS implementations;
  • We provide extensive evaluations of open-source and closed-source LLMs, showing that retrieval and fine-tuning on HLStrans significantly boost synthesis success rates and performance. By positioning HLStrans as both a resource and a benchmark, we aim to catalyze progress in LLM-powered hardware design and accelerate the integration of AI into future EDA workflows.
2 Background and Related works
@@
LLM aided C to HLS. There is an increasing body of literature on applying LLMs to generate HLS design from original C code. Collini et al. (2024) evaluates the basic task of translating naive C++ into synthesizable HLS C++. Bhattacharyya et al. (2024) demonstrates that LLMs can automate HLS pragmas and optimizations to produce synthesizable, high-performance RTL from C on image-processing benchmarks. Xu et al. (2024) presents an LLM-driven HLS program-repair framework that transforms C/C++ into synthesizable HLS-C. Xiong et al. (2024) extends this approach with retrieval-augmented generation and chain-of-thought prompting to deliver optimized HLS implementations across nine applications. However, to date, no work has evaluated LLM’s capabilities transforming C code to HLS codes on a large-scale dataset.
HLS code dataset. HLSsyn (Bai et al., 2023) focuses on incorporating a diverse set of optimization pragmas but contains only 42 kernels for training and evaluating design-quality prediction models. HLSDataset (Wei et al., 2023), which aggregates 34 data sources into roughly 18K samples, targets power, resource, and timing estimation. MLSBench (Goswami et al., 2022) is an open-source corpus produced with the Xilinx Vivado HLS flow; it covers 17 C/C++ and 13 SystemC benchmarks, but provides only HLS log files and reports. DB4HLS (Ferretti et al., 2021) introduced a database of more than 100,000 HLS design points generated from MachSuite via exhaustive design-space exploration. Likewise, Dai et al. (2018) released about 1,300 designs created from benchmarks. Despite these valuable resources, they suffer from three key limitations when used to evaluate LLMs’ ability to translate C code into HLS:
First, prior HLS datasets have primarily targeted quality-of-results (QoR) estimation rather than C-to-HLS code generation, and the underlying program sources are limited. Although varying tool configurations can yield many synthesized samples, the scarcity of distinct source programs prevents an LLM from learning diverse program structures needed for C-to-HLS tasks. Moreover, the selected programs are typically short, making them inadequate for fully evaluating an LLM’s capability.
Second, Existing datasets inadequately capture comprehensive C-to-HLS transformations, focusing largely on pragma insertion. Generating high-performance HLS code from standard C/C++ for FPGAs requires a series of systematic transformations to expose parallelism, optimize data movement, and conform to HLS-friendly coding styles. While the detailed transformations are in Appendix A.1, these transformations fall into five broad categories, shown in Figure 1.
Figure 1: C/C++ to HLS code transformation examples. T1: Apply loop tiling and local buffering to improve data locality. T2: Unroll inner loops to increase parallelism and throughput. T3: Convert floating-point to fixed-point types to reduce resource use and latency. T4: Replace standard math calls with HLS intrinsics (e.g. hls::sqrt) for synthesizable implementations. T5: Eliminate recursion by refactoring to iterative code so the design can be synthesized.
T1: Code Restructuring. Refactor algorithms to expose pipelining and dataflow, apply loop tiling, memory coalescing, ping-pong buffering, and reorganize control logic for parallel or streaming execution. T2: Directive (Pragma) Insertion. Place HLS pragmas to guide the tool scheduler, such as data flow, pipeline, loop partition, and interface specifications, to fine-tune performance and resource usage.T3: Data-Type Adaptation. Replace generic C types with HLS-specific arbitrary-precision types: convert floating point to fixed point (ap_fixed) for resource optimization, standard integers to bit-accurate (ap_int/ap_uint), and customize bit widths to match application precision requirements. T4: Transformation of Functions. Transform standard C functions into HLS-optimized kernels or intrinsics (such as converting the std::sqrt function to the hls::sqrt function) to better leverage FPGA fabric and specialized accelerators. T5: HLS-Compliant Coding Style. Eliminate unsupported C constructs such as dynamic memory allocation (malloc/free), recursion, and certain pointer arithmetic patterns; restructure code to use static arrays, simple loops, and explicit handshaking for communication.
Third, they are not organized as paired C-and-HLS examples and omit the corresponding testbenches needed for LLM-based HLS code optimization, which are not ready for LLM to verify its output.
@@
Table 1: Comparison of HLS datasets. QoR: quality of result prediction. Transformation: C to HLS transformations mentioned in 1. ✓: included. ✗: not included
Attributes Dai MLSBench DB4HLS HLSdataset HLSsyn HLStrans
Samples 1,300 6,000 124,106 18,876 42,000 124,200
Programs 65 30 19 34 42 309
Purpose QoR QoR QoR QoR QoR Code generation
Transformations T2 T2 T2 T1,T2 T2 T1, T2, T3, T4, T5
Testbench No No No No No Yes
Programs
CHStone(Hara et al., 2008)
Polybench(Pouchet & Yuki, 2012)
Rodinia(Che et al., 2009)
Machsuite(Reagen et al., 2014)
Rosetta(Zhou et al., 2018)
C2HLS(Collini et al., 2024)
PP4FPGA(Kastner et al., 2018)
Forgebench(Wanna et al., 2025)
HLSfactory(Abi-Karam et al., 2024)
Others (GitHub)
3 HLStrans Datasets Construction
Open-source HLS datasets are scarce and poorly structured, which limits their usefulness for training LLMs. We propose an automated pipeline to generate high-quality HLS datasets from existing resources. The pipeline has three stages: (1) collect high-quality human optimized open-source HLS examples; (2) perform targeted data augmentation on human optimized kernels to produce many viable candidates; (3) select the efficient HLS implementations from those candidates. Figure 2 shows our dataset construction process.
Figure 2: HLStrans Dataset Construction Process.
@@ 3.1 Dataset Collection
Firstly, we harvest code from GitHub, selecting repositories with at least ten stars. However, manually optimized codebases often exhibit inconsistent formatting and sparse documentation, which hinders LLM-driven code generation. Public kernels also frequently depend on unexpanded macros and bundle extraneous utility functions that obscure the core algorithm. To make C to HLS tasks readily consumable by LLMs, we package each design with the following files:
@@
  • Self-contained C++ testbench $tb$ includes all test cases and validation logic necessary to verify the kernel’s outputs against expected results. We manually write all the testbenches and adjust the optimized HLS code to ensure it passes all tests.
Therefore, we construct triples $(x,y,tb)$. If the original file $x$ is synthesisable, the execution cycles from synthesis reports of $y$ must be less than $y$. If the original file $x$ is not synthesisable, $y$ should be synthesisable.
@@ 3.2 Dataset Augmentation
Relying solely on collected repositories is insufficient because high-quality hardware codes are far scarcer than general software. To generate richer, more useful examples, we designed an automated dataset-augmentation framework that synthesizes additional C to HLS variants.
We formulate the C to HLS translation as a combinatorial search problem: selecting appropriate combinations of code transformations to meet performance and resource targets. Our approach proceeds in two stages. First, an LLM agent guided by Monte Carlo Tree Search (MCTS) proposes and explores semantic-preserving code transformations that expose parallelism and produce HLS-friendly structure. Second, for each candidate design we apply automated design-space-exploration (DSE) tools to tune pragmas and low-level implementation choices. Both stages are evaluated in the loop using EDA feedback (performance, resource utilization, and compilation outcomes), enabling MCTS and DSE to efficiently navigate the large, combinatorial action space (see Figure 3).
First, MCTS performs structured exploration by balancing the exploitation of high-reward actions with the exploration of uncertain or less-visited regions of the search space. The optimization policy is generated by the retrieval augmentation generation (RAG) module. The search is guided to choose the suitable policy by both the verification pipeline and a reward model. The reward model incorporates detailed feedback from the HLS toolchain, including synthesis success or failure, compile warnings, and performance metrics such as resource usage, latency, and throughput. This heuristic-driven strategy enables the agent to iteratively refine transformation sequences and produce more high-quality, synthesizable HLS designs. Second, HLS directive design space exploration using genetic algorithms (Ferikoglou et al., 2023) is adopted. It inserts pipeline, unroll, and partition pragmas to produce more effective data samples. Through iterative refinement, the framework converges toward optimized and synthesizable HLS code.
Figure 3: HLStrans Dataset Augmentation Framework.